OpenStudy (anonymous):

it take s70 oz of seed to feed 3500ft how much oz of seed would be needed to cover 5000ft

hero (hero):

Set up a rational proportion

OpenStudy (anonymous):

70/3500 =?/5000

hero (hero):

70/x=3500/5000

hero (hero):

I will explain in more detail in a few

OpenStudy (akashdeepdeb):

It is actually the same thing. And both equally make sense.

hero (hero):

I know but units should be accounted for as well

hero (hero):